OpenStudy (anonymous):

What would the taylor series centered at 0 be for the function below?

OpenStudy (anonymous):

\[f(x) = 2/(1-x^2)^3\]

OpenStudy (abb0t):

Taylor series: \[f(x) =\sum_{n=0}^{\infty} \frac{ f^{(n)}(a) }{ n! }(x-a)^n = f(x)+f'(a)(x-a)+\frac{ f'(a) }{ 2! }(x-a)^2+...\]

OpenStudy (abb0t):

taylor series is centered around x=0 I think

OpenStudy (anonymous):

I understand that but it's not looking for the expansion, it's looking for the actual series representation

OpenStudy (anonymous):

Your given function kind of resembles a geometric series. I'm sure there's some way you can use that as a starting point, but I'm not sure where to begin.

OpenStudy (anonymous):

WA gives an interesting looking result: http://www.wolframalpha.com/input/?i=power+series+2%2F%281-x%5E2%29%5E3

OpenStudy (anonymous):

Yeah, I looked at that and had no clue what to do with that, it didn't seem right but then again it is wolfram alpha